OpenStudy (anonymous):

CONVERGENT OR DIVERGENT? This is convergent (checked on Wolf), but I am having trouble showing it:

OpenStudy (anonymous):

\[\int\limits\limits_{\pi/2}^{\pi} 34\csc(x)dx\]

OpenStudy (anonymous):

@EarthCitizen @experimentX @apoorvk @FoolforMath @amistre64 @Mertsj

OpenStudy (anonymous):

\[\int\limits\limits\limits_{\pi/2}^{\pi} 34\csc(x)dx = \lim_{a \rightarrow \pi-} \int\limits\limits\limits_{\pi/2}^{a} 34\csc(x)dx\]

OpenStudy (anonymous):

\[\int\limits\limits_{}^{}\csc(x) dx = \ln|\csc(x) - \cot(x)| + C\] ^indefinite integral of csc(x), lets get that straight..

OpenStudy (experimentx):

singularity at pi/2 \[ \lim_{x->\pi} \ln|\csc x - \cot x| - \ln|\csc \pi/2 - \cot \pi/2|\] Doesn't it look like \( \infty - \infty \) try multiplying by conjugate and use L'Hospital rule

OpenStudy (experimentx):

Not sure though ... let me check from wolfram!!

OpenStudy (experimentx):

should be closed http://www.wolframalpha.com/input/?i=integrate+csc%28x%29+from+pi%2F2+to+pi

OpenStudy (experimentx):

Try to find the value of inner limit \[ \ln \left | \lim_{x\rightarrow \pi}\frac{\csc x - \cot x}{\csc (\pi/2) - \cot (\pi/2)}\right |\]

OpenStudy (experimentx):

Sorry that didn't work!!

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@amistre64. We have got up to the limit part, and there we are having trouble evaluating it. It HAS to be convergable. We checked on Wolf

OpenStudy (amistre64):

convergent or divergent would imply some trouble spots that need to be addresses

OpenStudy (amistre64):

1/sin is troublesome at sin = 0 when does sin = 0 on your interval

OpenStudy (amistre64):

and i see youve already determined the integral part ln(csc - cot) right?

OpenStudy (anonymous):

Yes.

OpenStudy (amistre64):

and the tricky part is the -F(a) condition eh

OpenStudy (anonymous):

Yea! If ONLY it were the OTHER way around

OpenStudy (anonymous):

Then I could exponentiate the ln's argument by -1.

OpenStudy (amistre64):

\[\lim_{a\to\ pi}-\ln (\csc a - \cot a)\]

OpenStudy (anonymous):

How???

OpenStudy (anonymous):

Where did that -1 come from?

OpenStudy (amistre64):

F(b) - F(a) is the solution

OpenStudy (amistre64):

you know F(b)

OpenStudy (anonymous):

But pi is the upper limit. How can you get the -1 in front of the ln?

OpenStudy (amistre64):

it is? such tiny little text

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):